An elevator has mass 650 kg , not including passengers. The elevator is designed to ascend, at constant speed, a vertical distance of 19.0 m (five floors) in 15.5 s , and it is driven by a motor that can provide up to 44 hp to the elevator.

Part A

What is the maximum number of passengers that can ride in the elevator? Assume that an average passenger has mass 65.0 kg.

Express your answer in integer

n = _______
Answer #1


power P = Eenrgy/time

P = mgh/t

1 hp = 746 Watts

Power Max 44 *746 Watts

speed of elevator = 19/15.5 = 1.22 m/s

Power = Force * Velocity

total weight = (650 + 65x) * 9.8

herr x is no. of passenegrs

Power for lifting = Force * velocity

P max = (650 + 65x) * 9.8*1.22 = 44* 746

solving for x

x = 32 passengers